# OMG: Towards Open-vocabulary Motion Generation via Mixture of Controllers

### 18 Dec 2023 v2

@@ 1 Introduction
Benefited from the existing text-annotated motion datasets [18, 47], recent advances generate diverse motions from text prompts, equipped with various generative models like VAEs [56, 76], autoregressive models [87, 32], and diffusion models [77, 89, 12, 11]. However, these methods heavily rely on the paired text-motion data with limited text annotations, and hence fall short of generalizing to unseen open-vocabulary text inputs. Imagine generating “Gollum breakdance footwork”, it requires out-of-domain generation ability with rich human knowledge to understand the motion traits in various words like characters (Gollum, ninja, etc.) or skills (breakdance, spinkick, etc.). To tackle the open-vocabulary texts, recent works [27, 41, 8] utilize the zero-shot text-image alignment ability of the large-scale pre-trained models, e.g., CLIP [57]. They adopt text-pose alignments and hence remove the reliance on pair-wise text and continuous motions. Yet, the discrete poses are insufficient to represent the rich motion traits, leading to unrealistic motion generation. On the other hand, recent advances [63, 88] enable impressive and controllable text-to-image generation, even from open-vocabulary text inputs. Revisiting their huge success, both the pretrain-then-finetune paradigm as well as scaling up the model have played an important role. Yet, adopting these strategies for text-to-motion generation is challenging. First, there exists a huge imbalance between the quantities of unpaired motion data and paired text-motion data. Second, for open-vocabulary texts, various tokens correspond to various motion traits, constituting a complex many-to-many problem.
@@ 2 Related Work
Conditional Motion Synthesis. Being able to generate realistic and contextually relevant motion sequences based on various types of conditions, conditional motion synthesis has received increasing attention in the field of motion generation. Common types of conditions include text [43, 3, 4, 76, 56, 84, 11, 19, 33, 66, 40], action [17, 55], music [34, 2, 38, 36], speech [20, 5, 6], control signals [71, 73, 54], action labels [17, 55, 81], incomplete motion sequences [13, 21], images [60, 10, 24, 39, 61] and scene [91]. The advent of Diffusion Models [25, 70] has given a big boost to text-driven motion synthesis. Kim et al. [33] develops a transformer-based diffusion model which could generate and edit motions well aligned with the given text. Motion Diffusion Model (MDM) [77] combines insights already well established in the motion generation domain and in each diffusion step predicts the sample rather than the noise. Motion Latent-based Diffusion model (MLD) [11] performs a diffusion process on the motion latent space. Besides the diffusion model, T2M-GPT [87] investigates a framework that combines VQ-VAE [79] and autoregressive model. However, since these models are trained only based on paired text-motion datasets such as HumanML3D [19], they cannot well handle unseen text prompts.
@@
Mixture-of-Experts. In the deep learning field, Mixture-of-Experts (MoE) [29, 67, 15] in a neural network architecture divides specific model parameters into distinct groups, referred to as “experts”. Eigen et al. [14] use a different gating network at each layer in a multi-layer network, increasing the number of effective experts by introducing an exponential number of paths through different combinations of experts at each layer. Other research that regards the entire model as an expert [37] also achieves good results. Moreover, endeavors are underway to enhance the training and inference speed within the MoE framework [23, 35]. The MoE paradigm is applicable not only to language processing tasks but also extends to visual models [62], Mixture-of-Modality-Experts in multi-modal transformers [68], as well as motion synthesis. Holden et al. [26] develop phase-functioned experts blended by pre-defined blending weights to control character at a specific phase, and Starke et al. [73, 72], Zhang et al. [85] uses a gating network to predict the blending weights, achieving impressive results of character control. Inspired by them, we devise text-token-specific experts that control the sub-motion for each text token.
@@ 3.1 Motion Diffusion Pre-training
In our experiments, $\mathcal{\lambda}_{vel}=30$ and $\mathcal{\lambda}_{foot}=30$. The number of diffusion timesteps $T$ is set to 1,000 with cosine noise level schedule [50]. We train all sizes of the model for 1M iterations with a batch size of $256$ using the AdamW [46] optimizer with a weight decay of 00, a maximum learning rate of $10^{-4}$, and a cosine LR schedule with 10K linear warmup steps. The models are trained using Pytorch with ZeRO [58] memory redundancy strategy on 8 NVIDIA A100 GPUs, and the largest model takes 1 week to train.
@@
Figure 4: Qualitative results generated by our model given various text prompts. Our model effectively captures the motion characteristics from either a single phrase or longer natural sentences.
Sliding Random Windows. During training, we propose sliding random windows that iterate over each motion state frame $x_{i}$ in the motion dataset to sample motion sequences starting from $x_{i}$. Specifically, for the $i^{th}$ frame, we randomly crop a sub-sequence $\mathbf{x}$ of length $l$ with a random window $[i,i+l)$, where $l\sim\mathcal{U}[1,L]$ is a uniform variable, and $L$ is the hyper-parameter that determines the maximum length of the generated motion sequence of our model. This pushes the model to learn the relations among both temporal frames and spatial features of a single motion state to process arbitrary lengths of motions even a single keyframe, to facilitate the following fine-grained control of sub-motions of arbitrary lengths. In our experiments, for balancing resolution and duration, the maximum length $L$ of sliding random windows is set to $300$ with the motion frame rate $30$.
@@ 3.2 Motion ControlNet
Training and Inference. The conditional denoiser $\mathcal{D}_{c}$ is also trained using the same objective $\mathcal{L}$ (Eq. 2) with trainable parameters in motion ControlNet. In our experiments, we employ the frozen CLIP-VIT-L/14 [57] text encoder to extract text embeddings at the final layer normalization with dimension $d_{c}=768$ and we truncate the text tokens with the maximum token number $77$. We set the down-projection latent dimension $d_{m}$ of MoC blocks to $256$, the expert pool size $K=12$, and attention mask sharpness $\gamma=24$ and threshold $\beta=0.25$. We train all the motion ControlNets for $500$ epochs with a batch size of $64$ using the AdamW optimizer with a weight decay of $51e-$ 5, a maximum learning rate of $3\times 10^{-5}$, and a cosine LR schedule with 1K linear warmup steps. In the training process, we randomly replace $\langle eos\rangle$ token with the empty token with 50% probability to increase the ability to capture global semantics in text tokens as a replacement.
@@ 4 Experiments
In this section, we first show our qualitative results in Fig. 4. Our method enables fine-grained control of complex and abstract motion trait descriptions. Then, we introduce dataset settings, and evaluation metrics (Sec. 4.1). We further compare our approach with various state-of-the-art methods on both in-domain and out-domain datasets (Sec. 4.2) and conduct extensive ablation studies on our model architecture (Sec. 4.3). We refer the reader to appreciate more qualitative results and details provided in supplements.
@@ 4.2 Comparison
We compare our approach with various state-of-the-art methods. Specifically, we apply conditional motion diffusion models, including MDM [77], MLD [11], MotionDiffuse [89], VAE-based model MotionCLIP [76], and auto-regressive model T2M-GPT [87]. Besides, we also apply MAA [8] based on the text-pose alignment method, and MotionGPT [32] utilizing motion-language pre-training. Some results on HumanML3D are borrowed from their own paper.
The quantitative results are presented in Tab. 2. Our method demonstrates the best text-to-motion alignment (R-Precision) in the in-domain evaluation. Moreover, among all diffusion-based methods, our model achieves the best FID. In terms of zero-shot performance, our approach achieves the best FID and CLIP-score, outperforming previous methods. This suggests a superior capability for high-quality motion generation and effective matching with zero-shot text prompts. As depicted in Fig. 5 of the qualitative results, our model demonstrates the capability to generate motions that show more realism, and better alignment with each motion characteristic description, whether in the sentence or phrase.

### 19 Mar 2024 v3

@@ 1 Introduction
Benefited from the existing text-annotated motion datasets [19, 49], recent advances generate diverse motions from text prompts, equipped with various generative models like VAEs [58, 78], autoregressive models [90, 34], and diffusion models [79, 92, 13, 12]. However, these methods heavily rely on the paired text-motion data with limited text annotations, and hence fall short of generalizing to unseen open-vocabulary text inputs. Imagine generating “Gollum breakdance footwork”, it requires out-of-domain generation ability with rich human knowledge to understand the motion traits in various words like characters (Gollum, ninja, etc.) or skills (breakdance, spinkick, etc.). To tackle the open-vocabulary texts, recent works [29, 43, 8] utilize the zero-shot text-image alignment ability of the large-scale pre-trained models, e.g., CLIP [59]. They adopt text-pose alignments and hence remove the reliance on pair-wise text and continuous motions. Yet, the discrete poses are insufficient to represent the rich motion traits, leading to unrealistic motion generation. On the other hand, recent advances [65, 91, 21, 9] enable impressive and controllable text-driven image and even video generation, even from open-vocabulary text inputs. Revisiting their huge success, both the pretrain-then-finetune paradigm as well as scaling up the model have played an important role. Yet, adopting these strategies for text-to-motion generation is challenging. First, there exists a huge imbalance between the quantities of unpaired motion data and paired text-motion data. Second, for open-vocabulary texts, various tokens correspond to various motion traits, constituting a complex many-to-many problem.
@@ 2 Related Work
Conditional Motion Synthesis. Being able to generate realistic and contextually relevant motion sequences based on various types of conditions, conditional motion synthesis has received increasing attention in the field of motion generation. Common types of conditions include text [45, 3, 4, 78, 58, 87, 12, 20, 35, 68, 42], action [18, 57], music [36, 2, 40, 38], speech [22, 5, 6, 95], control signals [73, 75, 56, 83], action labels [18, 57, 84], incomplete motion sequences [14, 23], images [62, 11, 26, 41, 63] and scene [94]. The advent of Diffusion Models [27, 72] has given a big boost to text-driven motion synthesis. Kim et al. [35] develops a transformer-based diffusion model which could generate and edit motions well aligned with the given text. Motion Diffusion Model (MDM) [79] combines insights already well established in the motion generation domain and in each diffusion step predicts the sample rather than the noise. Motion Latent-based Diffusion model (MLD) [12] performs a diffusion process on the motion latent space. Besides the diffusion model, T2M-GPT [90] investigates a framework that combines VQ-VAE [81] and autoregressive model. However, since these models are trained only based on paired text-motion datasets such as HumanML3D [20], they cannot well handle unseen text prompts.
@@
Mixture-of-Experts. In the deep learning field, Mixture-of-Experts (MoE) [31, 69, 16] in a neural network architecture divides specific model parameters into distinct groups, referred to as “experts”. Eigen et al. [15] use a different gating network at each layer in a multi-layer network, increasing the number of effective experts by introducing an exponential number of paths through different combinations of experts at each layer. Other research that regards the entire model as an expert [39] also achieves good results. Moreover, endeavors are underway to enhance the training and inference speed within the MoE framework [25, 37]. The MoE paradigm is applicable not only to language processing tasks but also extends to visual models [64], Mixture-of-Modality-Experts in multi-modal transformers [70], as well as motion synthesis. Holden et al. [28] develop phase-functioned experts blended by pre-defined blending weights to control character at a specific phase, and Starke et al. [75, 74], Zhang et al. [88] uses a gating network to predict the blending weights, achieving impressive results of character control. Inspired by them, we devise text-token-specific experts to control the corresponding sub-motions.
@@ 3.1 Motion Diffusion Pre-training
In our experiments, $\mathcal{\lambda}_{vel}=30$ and $\mathcal{\lambda}_{foot}=30$. The number of diffusion timesteps $T$ is set to 1,000 with cosine noise level schedule [52]. We train all sizes of the model for 1M iterations with a batch size of $256$ using the AdamW [48] optimizer with a weight decay of 00, a maximum learning rate of $10^{-4}$, and a cosine LR schedule with 10K linear warmup steps. The models are trained using Pytorch with ZeRO [60] memory redundancy strategy on 8 NVIDIA A100 GPUs. The largest model OMG-G takes 1500 GPU hours for pre-training.
@@
Figure 4: Qualitative results generated by our model given various unseen text prompts. Our model effectively captures the motion characteristics from either a single phrase or longer natural sentences.
Sliding Random Windows. During training, we propose sliding random windows that iterate over each motion state frame $x_{i}$ in the motion dataset to sample motion sequences starting from $x_{i}$. Specifically, for the $i^{th}$ frame, we randomly crop a sub-sequence $\mathbf{x}$ of length $l$ with a random window $[i,i+l)$, where $l\sim\mathcal{U}[1,L]$ is a uniform variable, and $L$ is the hyper-parameter that determines the maximum length of the generated motion sequence of our model. This pushes the model to learn the relations among both temporal frames and spatial features of a single motion state to process arbitrary lengths of motions even a single keyframe, to facilitate the following fine-grained control of sub-motions of arbitrary lengths. In our experiments, for balancing resolution and duration, the maximum length $L$ of sliding random windows is set to $300$ with the framerate set to $30$.
@@ 3.2 Motion ControlNet
Training and Inference. The conditional denoiser $\mathcal{D}_{c}$ is also trained using the same objective $\mathcal{L}$ (Eq. 2) with trainable parameters in motion ControlNet. In our experiments, we employ the frozen CLIP-VIT-L/14 [59] text encoder to extract text embeddings at the final layer normalization with dimension $d_{c}=768$ and we truncate the text tokens with the maximum token number $77$. We set the down-projection latent dimension $d_{m}$ of MoC blocks to $256$, the expert pool size $K=12$, and attention mask sharpness $\gamma=24$ and threshold $\beta=0.25$. We train all the motion ControlNets for $500$ epochs with a batch size of $64$ using the AdamW optimizer with a weight decay of $1\times 10^{-5}$, a maximum learning rate of $3\times 10^{-5}$, and a cosine LR schedule with 1K linear warmup steps. In the training process, we randomly replace $\langle eos\rangle$ token with the empty token with 50% probability to increase the ability to capture global semantics in text tokens as a replacement.
@@ 4 Experiments
We show qualitative results in Fig. 4. Our method enables fine-grained control of complex and abstract motion trait descriptions. We encourage the reader to appreciate more qualitative results provided in supplements and the video.
@@ 4.2 Comparison
We compare our approach with various state-of-the-art methods. Specifically, we apply conditional motion diffusion models, including MDM [79], MLD [12], MotionDiffuse [92], VAE-based model MotionCLIP [78], and auto-regressive model T2M-GPT [90]. Besides, we also apply MAA [8] based on the text-pose alignment method, and MotionGPT [34] utilizing motion-language pre-training. Some results on HumanML3D are borrowed from their own paper. The quantitative results are presented in Tab. 2. Our method demonstrates the best text-to-motion alignment (R-Precision) in the in-domain evaluation. Moreover, among all diffusion-based methods, our model achieves the best FID. In terms of zero-shot performance, our approach achieves the best FID and CLIP-score, outperforming previous methods. This suggests a superior capability for high-quality motion generation and effective matching with zero-shot text prompts. As depicted in Fig. 5 of the qualitative results, our model demonstrates the capability to generate motions that show more realism, and better alignment with each motion characteristic description, whether in the sentence or phrase.
@@ 5 Conclusion
Acknowledgement
This work was supported by National Key R&D Program of China (2022YFF0902301), Shanghai Local college capacity building program (22010502800). We also acknowledge support from Shanghai Frontiers Science Center of Human-centered Artificial Intelligence (ShangHAI).